During one month, a homeowner used 200 units of electricity and 130 units of gas for a total cost of $71.50. The next month, 220 units of electricity and 200 units of gas were used for a total cost of $87.20. Find the cost per unit of gas.

Answer by macston(5194)

e=cost per unit of electricity; g=cost per unit of gas
200e+130g=$71.50 Multiply each side by 20/13
(4000/13)e+200g=$110.00
220e+200g=$87.20 Subtract from above.
(4000/13)e-220e=$22.80
(4000/13)e-(2860/13)e=$22.80
(1140/13)e=$22.80
87.69e=$22.80
e=$0.26
200e+130g=$71.50
200($.26)+130g=$71.50
$52+130g=$71.50 Subtract $52 from each side
130g=$19.50 Divide each side by 130
g=$0.15 ANSWER: The cost of gas is $0.15 per unit.
CHECK:
220e+200g=$87.20
220($0.26)+200($0.15)=$87.20
$57.20+$30.00=$87.20
$87.20=$87.20